How do you figure this out when y is isolated on both equations:
y=3x+1
y=6x-1
Thank you

Found 3 solutions by scott8148, stanbon, orca:
Answer by scott8148(6628)

you substitute equals for equals

3x+1=6x-1 __ subtracting 3x __ 1=3x-1 __ adding 1 __ 2=3x __ dividing by 3 __ 2/3=x

substituting for x __ y=3(2/3)+1 __ y=3

Answer by stanbon(75887)

How do you figure this out when y is isolated on both equations:
y=3x+1
y=6x-1
------------
Since both are "y", they can be substituted to give:
3x + 1 = 6x -1
3x = 2
x = 2/3
------------------
Substitute into y=3x+1 to solve for "y":
y = 3(2/3) = 2
-----------------
Ans: x = 2/3 ; y = 2
==========================
Cheers,
Stan H.

Answer by orca(409)

The easiest way to solve this system of equations is:
As y is equal to both 3x + 1 and 6x - 1, so
6x - 1 = 3x + 1
Solving it for x, we have
3x = 2
so x = 2/3
Substituting x = 2/3 into y = 3x + 1, we have
y = 3*(2/3) + 1 = 3
*************************************************
ALTERNATIVE SOLUTION:
Change the system into standard form.
-3x + y = 1
-6x + y = -1
then solve it.